OpenStudy (braves12):

multiplying polynomials (a - b)(a 2 + ab + b 2)

OpenStudy (sleepyjess):

Is this supposed to be \((a-b)(a^2+ab+b^2)\)?

OpenStudy (braves12):

yes

OpenStudy (sleepyjess):

If so, then multiply the following... \(a*a^2\\a*ab\\a*b^2\\-b*a^2\\-b*ab\\-b*b^2\)

OpenStudy (anonymous):

a3+a2b+ab2 a2b+ab2+b3 vvvvvvvvvvvvv a3+a2b+ab2+b3

OpenStudy (solomonzelman):

this is one of the (a-b) to some power. To which, do you know?

OpenStudy (sleepyjess):

\(a*a^2=a^3\\a*ab=a^2b\\a*b^2=ab^2\\-b*a^2=-a^2b\\-b*ab=-ab^2\\-b*b^2=-b^3\\a^3+a^2b+ab^2-a^2b-ab^2-b^3\)

OpenStudy (braves12):

what about this one (a + 3)(a - 2)

OpenStudy (anonymous):

Mine is already completed and solved. And no not (a+3)(a-2)

OpenStudy (anonymous):

My example is pretty clear and my result is the same as you would get when multiplying polynomials

OpenStudy (sleepyjess):

Why do you have everything positive @AlondraAcosta ? You are multiplying 3 things by -b